OpenStudy (anonymous):

A ball is thrown from the ground at a velocity of 64 feet per second. The formula h = –16t2 + 64t gives the ball’s height, in feet, after t seconds. How many seconds will it take the ball to reach its maximum height?

OpenStudy (anonymous):

max is at the vertex, first coordinate of the vertex of a parabola (which is what you are looking for) is always \(-\frac{b}{2a}\) which in your case is \[-\frac{64}{2\times (-16)}\]

OpenStudy (anonymous):

better known as 2

OpenStudy (anonymous):

Can you answer this one? : What is the y-intercept of the function that passes through the points (–3, 12) and (4, –2)?